OpenStudy (anonymous):

integral of sec^4 x * cos^2x * tan ^7x dx tired and brain dead someone do it for me plz

OpenStudy (anonymous):

if you want an answer try http://www.wolframalpha.com/input/?i=sec^4+x+*+cos^2x+*+tan+^7x+dx

OpenStudy (anonymous):

pellet son thats a convenient cheater box u got there

OpenStudy (anonymous):

but you see it is simple answer and that is because \[\sec^4(x)\times \cos^2(x)=\frac{\cos^2(x)}{\cos^4(x)}=\sec^2(x)\]

OpenStudy (anonymous):

so really you just have \[\int\tan^7(x)\sec^2(x)dx\] and a simple u -sub gives it in one step

OpenStudy (anonymous):

oh wow thats embarrassing

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok one more

OpenStudy (anonymous):

right circular cylinder is inscribed in a cone with height h and base radius r Find the largest possible volume of such a cylinder